# Robust Score-Based Quickest Change Detection

### v3

@@ Abstract
Methods in the field of quickest change detection rapidly detect in real-time a change in the data-generating distribution of an online data stream. Existing methods have been able to detect this change point when the densities of the pre- and post-change distributions are known. Recent work has extended these results to the case where the pre- and post-change distributions are known only by their score functions. This work considers the case where the pre- and post-change score functions are known only to correspond to distributions in two disjoint sets. This work employs a pair of “least-favorable” distributions to robustify the existing score-based quickest change detection algorithm, the properties of which are studied. This paper calculates the least-favorable distributions for specific model classes and provides methods of estimating the least-favorable distributions for common constructions. Simulation results are provided demonstrating the performance of our robust change detection algorithm.
@@ I Introduction
This issue is partially addressed in [15] where the authors have proposed the SCUSUM algorithm, a Hyvärinen score-based ([16]) modification of the CUSUM algorithm for quickest change detection. It is shown in [15] that the SCUSUM algorithm is consistent and the authors also provide expressions for the average detection delay and the mean time to a false alarm. The Hyvärinen score is invariant to scale and hence can be applied to unnormalized models. This makes the SCUSUM algorithm highly efficient as compared to the classical CUSUM algorithm for high-dimensional models.
The main limitation of the SCUSUM algorithm is that its effectiveness is contingent on knowing the precise pre- and post-change unnormalized models, i.e., knowing the pre- and post-change models within a normalizing constant. In practice, due to a limited amount of training data, the models can only be learned within an uncertainty class. To detect the change effectively, an algorithm must be robust against these modeling uncertainties. The SCUSUM algorithm is not robust in this sense. Specifically, if not carefully designed, the SCUSUM algorithm can fail to detect several (in fact, infinitely many) post-change scenarios.
@@
In this paper, a robust score-based variant of the CUSUM algorithm, called RSCUSUM, is proposed for the problem of quickest change detection. Under the assumption that the pre- and post-change uncertainty classes are disjoint and convex, we show that the algorithm is robust, i.e., can consistently detect changes for every possible post-change model. This consistency is achieved by designing the RSCUSUM algorithm using the least favorable distributions from the pre- and post-change uncertainty classes.
The problem of optimal robust quickest change detection is studied in [17]. In a minimax setting, the optimal algorithm is the CUSUM algorithm designed using the least favorable distribution. The robust CUSUM test in [17] may suffer from two drawbacks: 1) It is a likelihood ratio-based test and hence may not be amenable to implementation in high-dimensional models. 2) The notion of least favorable distribution is defined using stochastic boundedness ([17]), which may be difficult to verify for high-dimensional data.
@@
    We propose a new robust score-based quickest change detection algorithm that can be applied to score-based models and unnormalized models, namely, statistical models whose density involves an unknown normalizing constant. We generalize the SCUSUM algorithm of [15] to propose a robust score-based variant which we refer to as RSCUSUM. As with the SCUSUM method, the role of Kullback-Leibler divergence in classical change detection is replaced with the Fisher divergence between the pre-and post-change distributions, and the role of log-likelihood ratios is replaced with differences between Hyvärinen scores ([16]). This algorithm is introduced in Section V.
@@
    Our generalized RSCUSUM algorithm can address unknown pre- and post-change models. Specifically, assuming that the pre- and post-change laws belong to known families of distributions which are convex, we identify a least favorable pair of distributions which are nearest in terms of Fisher divergence. We then show that the RSCUSUM algorithm can consistently detect each post-change distribution from each pre-change distribution from the relevant families, and is robust in this sense. This analysis is provided in Section VI.
@@
    We provide an effective method to identify the least favorable post-change distribution among pre- and post-change families. This is in contrast to the setup in [17] where a stochastic boundedness characterization makes it harder to identify the least favorable distribution. This result is presented in Section VII.
@@
    As with the SCUSUM algorithm, the RSCUSUM algorithm is a score-based algorithm where cumulative scores do not enjoy a standard martingale characterization. This differentiates RSCUSUM and SCUSUM from the classic CUSUM method, which leverages the fact that the likelihood ratios form a martingale under the pre-change model [9, 18]. We generalize the analysis of the delay and false-alarm of the SCUSUM method to the robust case. This analysis is presented in Section VI.
@@ II Problem Formulation
$\mathcal{L}_{\texttt{WADD}}(T)\overset{\Delta}{=}\sup_{\nu\geq 1}\text{ess}\sup\mathbb{E}_{\nu}[(T-\nu+1)^{+}|\mathcal{F}_{\nu-1}],$
where $(y)^{+}\overset{\Delta}{=}\max(y,0)$ for any $y\in\mathbb{R}$. Here $\text{ess}\sup$ is the essential supremum, i.e., the supremum outside a set of measure zero. We also consider the version of minimax metric introduced in [11], the worst conditional averaged detection delay (CADD):
$\mathcal{L}_{\texttt{CADD}}(T)\overset{\Delta}{=}\sup_{\nu\geq 1}\mathbb{E}_{\nu}[T-\nu|T\geq\nu].$
@@
$\text{ARL}(T)\overset{\Delta}{=}\mathbb{E}_{\infty}[T].$
We assume that pre- and post-change distributions are not precisely known. However, each is known within an uncertainty class. Let $\mathcal{G}_{\infty}$ and $\mathcal{G}_{1}$ be two disjoint classes of probability measures (We will discuss precise assumptions on them below). Then, we assume that the distributions $P_{\infty}$ and $P_{1}$ satisfy
$\begin{split}P_{\infty}&\in\mathcal{G}_{\infty}\\ P_{1}&\in\mathcal{G}_{1}.\end{split}$ (1)
@@
$\min_{T}\;\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathcal{L}_{\texttt{WADD}}(T)\;\quad\text{subject to}\;\quad\inf_{P_{\infty}\in\mathcal{G}_{\infty}}\mathbb{E}_{\infty}[T]\geq\gamma,$ (2)
where $\gamma$ is a constraint on the ARL. The delay $\mathcal{L}_{\texttt{WADD}}$ in the above problem is a function of the true pre- and post-change laws $P_{\infty},P_{1}$ and should be designated as $\mathcal{L}_{\texttt{WADD}}^{P_{\infty},P_{1}}.$ We will, however, suppress this notation and refer to $\mathcal{L}_{\texttt{WADD}}^{P_{\infty},P_{1}}$ by $\mathcal{L}_{\texttt{WADD}}$. Similarly, we will suppress the dependence of $\mathbb{E}_{\infty}[T]$ on the pre-change law $P_{\infty}$. Thus, the goal in this problem is to find a stopping time $T$ to minimize the worst-case detection delay, subject to a constraint $\gamma$ on $\mathbb{E}_{\infty}[T]$. We are also interested in the minimax formulation:
$\min_{T}\;\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathcal{L}_{\texttt{CADD}}(T)\;\quad\text{subject to}\;\quad\inf_{P_{\infty}\in\mathcal{G}_{\infty}}\mathbb{E}_{\infty}[T]\geq\gamma.$ (3)
Similarly, the delay $\mathcal{L}_{\text{CADD}}$ is a function of $P_{\infty},P_{1}$, but we suppress the notation and refer to $\mathcal{L}_{\text{CADD}}^{P_{\infty},P_{1}}$ as $\mathcal{L}_{\text{CADD}}$. For an overview of the literature on quickest change detection when the densities are not precisely known, we refer the readers to [1, 4, 2, 9, 7].
@@ III Optimal Solution Based on Likelihood Ratios
$\begin{split}&\Lambda(0)=0,\\ &\Lambda(n)\overset{\Delta}{=}\biggr{(}\Lambda(n-1)+\log\frac{p_{1}(X_{n})}{p_{\infty}(X_{n})}\biggr{)}^{+},\;\quad\forall n\geq 1,\end{split}$
@@
Here $\mathbb{D}_{\texttt{KL}}(P_{1}\|P_{\infty})$ is the Kullback-Leibler divergence between the post-change distribution and pre-change distribution:
@@
and the notation $g(c)\sim h(c)$ as $c\to c_{0}$ indicates that $\frac{g(c)}{h(c)}\to 1$ as $c\to c_{0}$ for any two functions $c\mapsto g(c)$ and $c\mapsto h(c)$.
If the post-change densities $p_{\infty},p_{1}$ are not known and are assumed to belong to families $\mathcal{G}_{\infty},\mathcal{G}_{1}$, then the test is designed using the least favorable distributions. Specifically, in [17], it is assumed that there are densities $\bar{p}_{\infty}\in\mathcal{G}_{\infty},\bar{p}_{1}\in\mathcal{G}_{1}$ such that for every $p_{\infty}\in\mathcal{G}_{\infty}$ and $p_{1}\in\mathcal{G}_{1}$,
@@
If such densities $\bar{p}_{\infty},\bar{p}_{1}$ exist in the pre- and post-change families, then the robust CUSUM is defined as the CUSUM test with $\bar{p}_{\infty},\bar{p}_{1}$ used as the pre- and post-change densities:
$\begin{split}&\bar{\Lambda}(0)=0,\\ &\bar{\Lambda}(n)\overset{\Delta}{=}\biggr{(}\bar{\Lambda}(n-1)+\log\frac{\bar{p}_{1}(X_{n})}{\bar{p}_{\infty}(X_{n})}\biggr{)}^{+},\forall n\geq 1.\end{split}$
Such a test is exactly optimal for the problem of [7] under additional assumptions on the smoothness of densities, and asymptotically optimal for the problem in [11]. We refer the reader to [17] for a more precise optimality statement.
We note that in the literature on quickest change detection, the issue of the unknown post-change model has also been addressed by using a generalized likelihood ratio (GLR) test or a mixture-based test. While these tests have strong optimality properties, they are computationally even more expensive than the robust test described above; see [7, 9, 4].
@@ IV Quickest Change Detection in Unnormalized and Score-Based Models
In a recent work [15], the authors have developed a score-based CUSUM (SCUSUM) algorithm to detect changes in unnormalized and score-based models and also obtained its performance characteristics. In the score-based theory in [15], the Kullback-Leibler divergence is replaced by the Fisher divergence (defined precisely below) between the pre- and post-change densities.
@@
$\mathcal{S}_{\texttt{H}}(x,P)\overset{\Delta}{=}\frac{1}{2}\left\|\nabla_{x}\log p(x)\right\|_{2}^{2}+\Delta_{x}\log p(x),$
whenever it can be well defined. We assume that the Hyvärinen scores is well defined for all $P\in\mathcal{G}_{\infty}\cup\mathcal{G}_{1}$ by Assumption IV.4. Here, $\|\cdot\|_{2}$ denotes the Euclidean norm, $\nabla_{x}$ and $\Delta_{x}=\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}^{2}}$ respectively denote the gradient and the Laplacian operators acting on $X=(x_{1},\cdots,x_{d})^{\top}$.
@@
$\mathbb{D}_{\texttt{F}}(P\|Q)\overset{\Delta}{=}\mathbb{E}_{X\sim P}\left[\frac{1}{2}\left\|\nabla_{{x}}\log p(X)-\nabla_{{x}}\log q(X)\right\|_{2}^{2}\right],$ (8)
@@
Clearly, $\nabla_{{x}}\log p(x)$, $\nabla_{{x}}\log q(x)$, and $\Delta_{x}\log q(x)$ remain invariant if $p$ and $q$ are scaled by any positive constant with respect to $x$. Hence, the Fisher divergence and the Hyvärinen Score remain scale-variant concerning an arbitrary constant scaling of density functions.
@@
$Y(n)\overset{\Delta}{=}(Y(n-1)+\mathcal{S}_{\texttt{H}}(X_{n},P_{\infty})-\mathcal{S}_{\texttt{H}}(X_{n},P_{1}))^{+},\;\forall n\geq 1.$
@@
  1. 1.
    The SCUSUM algorithm is consistent. Specifically, the drift of the statistic is negative before the change and is positive after the change:
@@
  2. 2.
    Let there exist a $\lambda>0$ such that
@@
    Thus, setting $\omega=\frac{\log\gamma}{\lambda}$ in (9) implies
@@
  3. 3.
    Finally, with $\omega=\frac{\log\gamma}{\lambda}$, the delay performance is given by

### v4

@@ Abstract
Methods in the field of quickest change detection rapidly detect in real-time a change in the data-generating distribution of an online data stream. Existing methods have been able to detect this change point when the densities of the pre- and post-change distributions are known. Recent work has extended these results to the case where the pre- and post-change distributions are known only by their score functions. This work considers the case where the pre- and post-change score functions are known only to correspond to distributions in two disjoint sets. This work selects a pair of least-favorable distributions from these sets to robustify the existing score-based quickest change detection algorithm, the properties of which are studied. This paper calculates the least-favorable distributions for specific model classes and provides methods of estimating the least-favorable distributions for common constructions. Simulation results are provided demonstrating the performance of our robust change detection algorithm.
@@ I Introduction
This issue is partially addressed in [15] which proposed the SCUSUM algorithm, a Hyvärinen score-based ([16]) modification of the CUSUM algorithm for quickest change detection. In [15] it is shown that the SCUSUM algorithm is consistent, and expressions for the average detection delay and the mean time to a false alarm of SCUSUM are provided. The Hyvärinen score is invariant to scale and hence can be applied to unnormalized models. This makes the SCUSUM algorithm highly efficient as compared to the classical CUSUM algorithm for high-dimensional models.
The main limitation of the SCUSUM algorithm is that its effectiveness is contingent on knowing the precise pre- and post-change unnormalized or score-based models, i.e., knowing the pre- and post-change models within a normalizing constant or knowing the gradient of log density. In practice, due to a limited amount of training data, the models can only be learned within an uncertainty class. To detect the change effectively, an algorithm must be robust against these modeling uncertainties. The SCUSUM algorithm is not robust in this sense. Specifically, if not carefully designed, the SCUSUM algorithm can fail to detect several (in fact, infinitely many) pre- and post-change scenarios.
@@
In this paper, a robust score-based variant of the CUSUM algorithm, called RSCUSUM, is proposed for the problem of quickest change detection. Under the assumption that the pre- and post-change uncertainty classes are disjoint and convex, we show that the algorithm is robust, i.e., can consistently detect changes for every possible pre- and post-change model. This consistency is achieved by designing the RSCUSUM algorithm using a pair of distributions from the pre- and post-change uncertainty classes that are least favorable in a well-defined sense.
The problem of optimal robust quickest change detection is studied in [17]. In a minimax setting, the optimal algorithm is the CUSUM algorithm designed using the least favorable distributions. The robust CUSUM test in [17] may suffer from two drawbacks: 1) It is a likelihood ratio-based test and hence may not be amenable to implementation in high-dimensional models. 2) The notion of least favorable distribution is defined using stochastic boundedness ([17]), which may be difficult to verify for high-dimensional data.
@@
    We propose a new robust score-based quickest change detection (RSCUSUM) algorithm that can be applied to score-based models and unnormalized models. As with the SCUSUM method from [15], the role of Kullback-Leibler divergence in classical change detection is replaced with the Fisher divergence between the pre- and post-change distributions, and the role of log-likelihood ratios is replaced with differences between Hyvärinen scores ([16]). This algorithm is introduced in Section V.
@@
    The RSCUSUM algorithm can address unknown pre- and post-change models. Specifically, assuming that the pre- and post-change laws belong to known families of distributions that are convex, we identify a least favorable pair of distributions that are nearest in terms of Fisher divergence. We then show that the RSCUSUM algorithm can consistently detect each post-change distribution from each pre-change distribution from the relevant families, and is robust in this sense. This analysis is provided in Section VI.
@@
    We provide an effective method to identify the least favorable distributions among pre- and post-change families. This is in contrast to the setup in [17], where a stochastic boundedness characterization makes it harder to identify the least favorable distributions. This result is presented in Section VII.
@@
    As with the SCUSUM algorithm, the RSCUSUM algorithm is a score-based algorithm where cumulative scores do not enjoy a standard martingale characterization. This differentiates the RSCUSUM algorithm and the SCUSUM algorithm from the classical CUSUM algorithm, which leverages the fact that the likelihood ratios form a martingale under the pre-change model [9, 18]. We generalize the analysis of the delay and false alarm of the SCUSUM method to the robust case. This analysis is presented in Section VI.
@@ II Problem Formulation
$\mathcal{L}_{\texttt{{WADD}}}(T)=\sup_{\nu\geq 1}\text{ess}\sup\mathbb{E}_{\nu}[(T-\nu+1)^{+}|\mathcal{F}_{\nu-1}],$
where $(y)^{+}=\max(y,0)$ for any $y\in\mathbb{R}$. Here $\text{ess}\sup$ is the essential supremum, i.e., the supremum outside a set of measure zero. We also consider the version of minimax metric introduced in [11], the worst conditional averaged detection delay (CADD):
$\mathcal{L}_{\texttt{{CADD}}}(T)=\sup_{\nu\geq 1}\mathbb{E}_{\nu}[T-\nu|T\geq\nu].$
@@
$\text{ARL}(T)=\mathbb{E}_{\infty}[T].$
We assume that pre- and post-change distributions are not precisely known. However, each is known within an uncertainty class. Let $\mathcal{G}_{\infty}$ and $\mathcal{G}_{1}$ be two disjoint classes of probability measures. Then, we assume that the distributions $P_{\infty}$ and $P_{1}$ satisfy
$\begin{split}P_{\infty}&\in\mathcal{G}_{\infty},\\ P_{1}&\in\mathcal{G}_{1}.\end{split}$ (1)
@@
$\min_{T}\;\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathcal{L}_{\texttt{{WADD}}}(T)\;\;\text{ s.t.}\;\inf_{P_{\infty}\in\mathcal{G}_{\infty}}\mathbb{E}_{\infty}[T]\geq\gamma,$ (2)
where $\gamma$ is a constraint on the ARL. Precise assumptions on the families $\mathcal{G}_{1}$ and $\mathcal{G}_{\infty}$ will be discussed below.
The delay $\mathcal{L}_{\texttt{{WADD}}}$ in the above problem is a function of the true pre- and post-change laws $P_{\infty},P_{1}$ and should be designated as $\mathcal{L}_{\texttt{{WADD}}}^{P_{\infty},P_{1}},$ but we suppress the notation and refer to this quantity simply as $\mathcal{L}_{\texttt{{WADD}}}$. Similarly, we will suppress the dependence of $\mathbb{E}_{\infty}[T]$ on the pre-change law $P_{\infty}$. Thus, the goal in this problem is to find a stopping time $T$ to minimize the worst-case detection delay, subject to a constraint $\gamma$ on $\mathbb{E}_{\infty}[T]$.
We are also interested in the minimax formulation:
$\min_{T}\;\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathcal{L}_{\texttt{{CADD}}}(T)\;\text{ s.t. }\;\inf_{P_{\infty}\in\mathcal{G}_{\infty}}\mathbb{E}_{\infty}[T]\geq\gamma.$ (3)
The delay $\mathcal{L}_{\text{CADD}}$ is also a function of $P_{\infty},P_{1}$, but we suppress the notation and refer to $\mathcal{L}_{\text{CADD}}^{P_{\infty},P_{1}}$ as $\mathcal{L}_{\text{CADD}}$. For an overview of the literature on quickest change detection when the densities are not precisely known, we refer the readers to [1, 4, 2, 9, 7].
@@ III Optimal Solution Based on Likelihood Ratios
$\begin{split}&\Lambda(0)=0,\\ &\Lambda(n)=\biggr{(}\Lambda(n-1)+\log\frac{p_{1}(X_{n})}{p_{\infty}(X_{n})}\biggr{)}^{+},\;\quad\forall n\geq 1,\end{split}$
@@
Here $\mathbb{D}_{\texttt{{KL}}}(P_{1}\|P_{\infty})$ is the Kullback-Leibler divergence between the post-change distribution and the pre-change distribution:
@@
and the notation $g(c)\sim h(c)$ as $c\to c_{0}$ indicates that ${g(c)}/{h(c)}\to 1$ as $c\to c_{0}$ for any two functions $c\mapsto g(c)$ and $c\mapsto h(c)$.
If the densities $p_{\infty},p_{1}$ are not known and are assumed to belong to families $\mathcal{G}_{\infty},\mathcal{G}_{1}$, i.e., the families $\mathcal{G}_{\infty},\mathcal{G}_{1}$ are not singleton sets, then the optimal test is designed using the least favorable distributions in a stochastic bounded sense. Specifically, in [17], it is assumed that there are densities $\bar{p}_{\infty}\in\mathcal{G}_{\infty},\bar{p}_{1}\in\mathcal{G}_{1}$ such that for every $p_{\infty}\in\mathcal{G}_{\infty}$ and $p_{1}\in\mathcal{G}_{1}$,
@@
If such densities $\bar{p}_{\infty},\bar{p}_{1}$ exist in the pre- and post-change families, then the optimal algorithm is the robust CUSUM algorithm designed with $\bar{p}_{\infty},\bar{p}_{1}$ used as the pre- and post-change densities:
$\bar{\Lambda}(0)=0,\;\quad\quad\bar{\Lambda}(n)=\biggr{(}\bar{\Lambda}(n-1)+\log\frac{\bar{p}_{1}(X_{n})}{\bar{p}_{\infty}(X_{n})}\biggr{)}^{+},\;\forall n\geq 1.$
The optimality in [17] is established under additional assumptions on the smoothness of densities. We refer the readers to [17] for a more precise optimality statement.
We note that in the literature on quickest change detection, the issue of the unknown post-change model has been addressed by using a generalized likelihood ratio (GLR) test or a mixture-based test. While these tests have strong optimality properties, they are computationally expensive to implement; see [7, 9, 4].
@@ IV Quickest Change Detection in Unnormalized and Score-Based Models
In [15], a score-based CUSUM (SCUSUM) algorithm was developed to detect changes in unnormalized and score-based models and performance characteristics of this algorithm were obtained. In the score-based theory in [15], the Kullback-Leibler divergence is replaced by the Fisher divergence (defined precisely below) between the pre- and post-change densities.
@@
$\mathcal{S}_{\texttt{{H}}}(x,P)=\frac{1}{2}\left\|\nabla_{x}\log p(x)\right\|_{2}^{2}+\Delta_{x}\log p(x),$
whenever it can be well defined, where
$\Delta_{x}=\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}^{2}}$
and $\nabla_{x}$ are the Laplacian and gradient operators acting on $X=(x_{1},\cdots,x_{d})^{\top}$, and where $\|\cdot\|_{2}$ denotes the Euclidean norm. We assume that the Hyvärinen score is well defined for all $P\in\mathcal{G}_{\infty}\cup\mathcal{G}_{1}$ by Assumption IV.4.
@@
$\mathbb{D}_{\texttt{{F}}}(P\|Q)=\mathbb{E}_{X\sim P}\left[\frac{1}{2}\left\|\nabla_{{x}}\log p(X)-\nabla_{{x}}\log q(X)\right\|_{2}^{2}\right],$ (8)
@@
Clearly, $\nabla_{{x}}\log p(x)$, $\nabla_{{x}}\log q(x)$, and $\Delta_{x}\log q(x)$ remain invariant if $p$ and $q$ are scaled by any positive constant with respect to $x$. Hence, the Fisher divergence and the Hyvärinen Score remain scale-invariant concerning an arbitrary constant scaling of density functions.
@@
$Y(n)=(Y(n-1)+\mathcal{S}_{\texttt{{H}}}(X_{n},P_{\infty})-\mathcal{S}_{\texttt{{H}}}(X_{n},P_{1}))^{+},\;\forall n\geq 1.$
@@
(a) Consistency: The SCUSUM algorithm is consistent. Specifically, the drift of the statistic is negative before the change and is positive after the change:
@@
(b) Average Run Length: Let there exist a $\lambda>0$ such that
@@
Thus, setting $\omega={(\log\gamma)}/{\lambda}$ in (9) implies
@@
(c) Expected Detection Delay: Finally, with $\omega={(\log\gamma)}/{\lambda}$, the delay performance is given by